The life of a 60- watt light bulb in hours is known to be normally distributed with σ = 25
hours. Create 5 different random samples of 100 bulbs each which has a mean life of x_bar ~
1000 hours and perform one-way ANOVA with state it.


In this problem how the value is getting ∑X ij =499712
∑Xij^2 =499691630

please explain this to me sir
1
Expert's answer
2020-08-20T18:12:01-0400

The total sample size is "N=500." Therefore, the total degrees of freedom are:



"df_{total}=500-1=499"

The between-groups degrees of freedom are "df_{between}=5-1=4," and the within-groups degrees of freedom are:



"df_{within}=df_{total}-df_{between}=499-4=495"

"\begin{matrix}
 Sum= & 99645 & 99684 & 99687 & 99699 & 99592 \\
 Average= & 996.45 & 996.84 & 996.87 & 996.99 & 995.92
\end{matrix}"


"\begin{matrix}
 \displaystyle\sum_{i}X_{ij}^2= \\
 99313355 & 99390488 & 99397147 & 99420471 & 99420471
\end{matrix}"


"\begin{matrix}
 \displaystyle St.Dev.= \\
 14.939 & 14.733 & 14.964 & 14.759 & 10.249
\end{matrix}"


"SS_i=(n_i-1)s_i^2"

"(100-1)14.939^2=22094.75"

"(100-1)14.733^2=21489.44"

"(100-1)14.964^2=22167.31"

"(100-1)14.759^2=21564.99"

"(100-1)10.249^2=10399.36"


"\begin{matrix}
 \displaystyle SS= \\
 22094.75 & 21489.44 & 22167.31 & 21564.99 & 10399.36
\end{matrix}"


"\begin{matrix}
 n= & 100 & 100 & 100 & 100 & 100 \\
\end{matrix}"


"\displaystyle\sum_{i,j}X_{ij}="

"=99645 +99684 + 99687 + 99699 + 99592="

"=498307"


"\displaystyle\sum_{i,j}X_{ij}^2="

"=99313355+ 99390488 + 99397147+99420471 + 99420471="

"=496717525"



"SS_{total}=\displaystyle\sum_{i,j}X_{ij}^2-\dfrac{1}{N}(\displaystyle\sum_{i,j}X_{ij})^2=97792.502""SS_{within}=22094.75 + 21489.44 +22167.31+""+21564.99+ 10399.36=97715.85""SS_{between}=\displaystyle\sum_{j}n_j(\bar{x}_j-\bar{\bar{x}})^2=76.652""\u200b""MS_{between}=\dfrac{SS_{between}}{df_{between}}=\dfrac{76.652}{4}=19.163""MS_{within}=\dfrac{SS_{within}}{df_{within}}=\dfrac{97715.85}{495}=197.406""F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{19.163}{197.406}=0.097"

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1=\mu_2=\mu_3=\mu_4=\mu_5"

"H_1:" Not all means are equal.

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df_1=4" and "df_2=4," therefore, the rejection region for this F-test is "R=\{F:F>F_c=2.39\}."

Test Statistics



"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{19.163}{197.406}=0.097"



Since it is observed that "F=0.097<2.39=F_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=0.9834," and since  "p=0.9834\geq0.05," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.